# Model-free source seeking of exponentially convergent unicycle: theoretical and robotic experimental results

### v2

@@ 4.2 Simulation Results
In this subsection, we compare the traditional ESC design commonly found in the literature that is based on first-order Lie bracket design (durr2013lie) with the proposed exponentially convergent unicycle ESC design developed for a fourth-order objective function based on third-order Lie bracket. The objective function used in this study is defined as
@@
Figure 3 illustrates the time histories of $x$ and $y$ as well as the planar trajectory of the unicycle. The results clearly demonstrate that the proposed exponentially convergent ESC achieves significantly faster convergence toward the desired equilibrium compared with the traditional ESC method from literature based on first-order Lie bracket design (durr2013lie). Specifically, for the fourth-order objective function considered, the traditional ESC fails to converge within the 100 s simulation window, whereas the proposed design successfully converges within approximately 20 s. Lastly, the proposed exponentially convergent unicycle ESC exhibits a higher radius of oscillation compared to the traditional ESC method. This behavior is due to the control law itself requiring a higher amplitude for the perturbation signal input as observed in grushkovskaya2025extremum; pokhrel2023higher. Techniques to attenuate oscillation such as those in pokhrel2023control; grushkovskaya2024step; yilmaz2025unbiased can possibly resolve this issue and can be applied to the exponentially convergent unicycle ESC. We also note that attenuation of oscillations in the unicycle-based ESC has been successfully tried in robotic experiments as in elgohary2025model.
@@
Figure 3: Simulation results comparing the proposed exponentially convergent unicycle ESC with the traditional ESC based on first-order Lie bracket design from durr2013lie for a fourth-order objective function: (a) $x$ position, (b) $y$ position, and (c) planar trajectory of the proposed unicycle design.
@@ 4.3 Experimental Results
We note here that our design parameters meet the condition of Theorem 1 (see Remark 1). The experimental results are presented in Figures 4 and 5. As shown in Figure 4, the traditional ESC method based on first-order Lie bracket from durr2013lie fails to converge to the true minimum of the objective function even after 1200 s, instead settling near (x,y)=(0.98,1.93)(x,y)=(0.98,-1.93). Due to the fourth-order objective function, the traditional ESC method will not be able to excite high-order Lie brackets and hence it is not possible to guarantee exponential convergence for this approach. Instead, the traditional ESC method will follow a polynomial decay rate, which will be much slower than an exponential decay rate as discussed and clarified in grushkovskaya2025extremum.
@@
In contrast, the exponentially convergent unicycle ESC design converges much faster, first reaching the minimum at approximately 350 s and remaining about the true extremum for the duration of the experiment. The new ESC design clearly outperforms the classic literature ESC design and is validated by its ability to reach the minimum in significantly less time than the traditional approach. It is also important to note that a fourth-order objective function exhibits relatively flat regions near the minimum compared to a second-order objective function. The exponentially convergent unicycle ESC design can successfully steer the robot through these flat regions, as evident in Figure 6. Moreover, the variations in the objective function values become very small when the robot is close to the extremum point, confirming that steady-state convergence has been achieved. It is also important to highlight that the proposed design performed well even with expected resolution issues from the motion capturing system feedback due to the very small values of the fourth-order objective function, especially near the extremum. This is a positive indication about the robustness of the proposed design. The reader is directed to watch the experiment in our YouTube channel (video_known_J).
@@
Figure 4: Experimental results for the traditional ESC design based on first-order Lie bracket from durr2013lie with a fourth-order objective function. (a) $x$ position (b) $y$ position (c) planar trajectory of the robot. Note that the x-position converges to $0.98$ meters (black-dashed), resulting in a 0.02 meter error, and the y-position converges to 1.93-1.93 (black-dashed), which results in a 0.07 meter error. The true minimum is in red.
@@
Figure 5: Experimental results for the exponentially convergent unicycle ESC design with a fourth-order objective function. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (video_known_J).
@@
The results of this experiment are presented in Figure 7. As shown, the robot successfully reaches and oscillates around the location of maximum light intensity. The light sensor measurements decrease as the robot approaches the source, and the mean value of the signal is minimized toward the end of the experiment, despite small oscillations caused by the robot’s continuous motion. These results further confirm that the proposed unicycle ESC design can autonomously steer the robot through an unknown and spatially varying light field, thereby validating its real-time, model-free source-seeking capability. The reader is directed to watch this experiment video in YouTube (video_light).
Figure 7: Experimental results for light source-seeking with the exponentially convergent ESC model. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (video_light).
@@ 4.6 Response to Different Initial Values in Simulations and Experiment
Furthermore, we extend the distance between the initial and desired positions and show, in simulation, the ability of the exponentially convergent unicycle ESC to reach the desired position from a further distance. The initial condition for this simulation is taken as, x(0),y(0)=0.3,0.7x(0),\ y(0)=-0.3,\ -0.7 and the result is shown in Figure 13. The unicycle is still able to converge to the desired position even when starting further away from the desired position.
Figure 13: Simulation result for an initial position that is further away from the desired position with the exponentially convergent unicycle ESC. x(0),y(0)=0.3,0.7x(0),\ y(0)=-0.3,\ -0.7.
In the experiment, due to the hardware limitations of the TurtleBot3 robot, we were not able to test initial conditions that were further away from the desired position.
Overall, the exponentially convergent unicycle ESC does not show dependence on the initial position directions with respect to the extremum in both simulations and experiments. However, a gap between theory/simulations on the one hand and experiments on the other, was observed as the exponentially convergent unicycle has hardware limitations preventing it from converging if the initial condition is set very far from the extremum; a solution for that could be a future work that develops a bounded update rate design upgrade for the exponentially convergent ESC, analogous to the bounded update design upgrade in scheinker2017bounded with respect to the traditional ESC design in durr2013lie.
@@ 4.7 Comparison between Simulation and Experiment Results
To begin, we note that the unicycle dynamics model does not perfectly capture the true dynamics of the TurtleBot3 robot. Inertial terms are not captured in the unicycle dynamics as the unicycle is treated as a point mass. Also, dissipation terms, such as friction, are not captured within the unicycle dynamics. For these reasons, it is not possible to directly equate the simulation and experiment parameters. That is, Table 1 and Table 2 are not aligned in terms of the parameters, partly due to the differences in the unicycle model and the real-world system of TurtleBot3. To compare the simulation results with its theoretically predicted behavior vs. the experimental result, we aimed to match the initial convergence rate for the states between the two systems (simulation-based and experiment-based) in a similar fashion to what could be found in ESC literature when one compares the convergence rate of two different ESC systems, see for example the comparisons in elgohary2025extremum. Analyzing Table 1 and Table 2, we see that at $t=0$ s, $\dot{y}(0)$ is the same for both systems and the only difference between the two systems is in $\dot{x}(0)$ with the simulation-based system possessing a much higher initial convergence rate in $\dot{x}$ compared to the experiment-based system. That is, for the experiment-based system, $\dot{x}=0.2199$, and for the simulation-based system, $\dot{x}(0)=705.724$. We now introduce a new set of parameters, in Table 4, to reduce the initial convergence rate of $\dot{x}$ in the simulation-based system as much as possible to allow for a realtively fairer comparison with the experiment-based system. The experimental parameters are the same as in Table 2. With the new parameters in Table 4 for the simulation-based system, the initial convergence rate for $\dot{x}$ becomes $\dot{x}(0)=5.5382$. Now, the results that compare the experiment-based system and the simulation-based system that theoretically is guaranteed to have exponential convergence rate are provided in Figure 14. We observe that the simulation-based system, even with relatively higher initial convergence rate in $\dot{x}$ (i.e., more advantageous initial speed), is slower than the experimental-based system. This shows that the proposed unicycle ESC follows the theoretical predictions in Section 3 and their simulations; however, one need to keep in mind the reported limitation in Subsection 4.6 where the real-world system may not converge for initial conditions that are very far from the extremum.
@@ 5 Conclusion and future work
This paper provided a novel, first-of-its-kind exponentially convergent unicycle design inspired by significant recent results in higher-order Lie bracket approximations (grushkovskaya2025extremum; pokhrel2023higher). We proved exponential stability of the proposed design for objective functions that behave locally as a fourth-degree polynomial. Additionally, we validated the results by simulations and, more importantly, via experiments. In simulations, we showed the effective response of the proposed design under different initial conditions and under added levels of measurement delays/noise. In our experiments, we did not only validate the ability of the proposed design to operate with known objective functions and with different initial conditions, but we also validated its ability to operate in a complete model-free condition in a light source-seeking experiment. Furthermore, we compared for a case study the experimental results behavior vs. a simulation result behavior to confirm that the experimental result follow the predicted theoretical behavior.
In the future, we aim at generalizing the results here in this paper to general higher-degree polynomial objective functions and expand the proposed design into different forms that may reduce oscillations and posses bounded update rate.

### v3

@@ 4.2 Simulation Results
In this subsection, we compare the traditional ESC design commonly found in the literature that is based on first-order Lie bracket design (Dürr et al. (2013)) with the proposed exponentially convergent unicycle ESC design developed for a fourth-order objective function based on third-order Lie bracket. The objective function used in this study is defined as
@@
Figure 3 illustrates the time histories of $x$ and $y$ as well as the planar trajectory of the unicycle. The results clearly demonstrate that the proposed exponentially convergent ESC achieves significantly faster convergence toward the desired equilibrium compared with the traditional ESC method from literature based on first-order Lie bracket design (Dürr et al. (2013)). Specifically, for the fourth-order objective function considered, the traditional ESC fails to converge within the 100 s simulation window, whereas the proposed design successfully converges within approximately 20 s. Lastly, the proposed exponentially convergent unicycle ESC exhibits a higher radius of oscillation compared to the traditional ESC method. This behavior is due to the control law itself requiring a higher amplitude for the perturbation signal input as observed in Grushkovskaya and Eisa (2025); Pokhrel and Eisa (2023b). Techniques to attenuate oscillation such as those in Pokhrel and Eisa (2023a); Grushkovskaya and Ebenbauer (2024); Yilmaz et al. (2025) can possibly resolve this issue and can be applied to the exponentially convergent unicycle ESC. We also note that attenuation of oscillations in the unicycle-based ESC has been successfully tried in robotic experiments as in Elgohary et al. (2025).
@@
Figure 3: Simulation results comparing the proposed exponentially convergent unicycle ESC with the traditional ESC based on first-order Lie bracket design from Dürr et al. (2013) for a fourth-order objective function: (a) $x$ position, (b) $y$ position, and (c) planar trajectory of the proposed unicycle design.
@@ 4.3 Experimental Results
We note here that our design parameters meet the condition of Theorem 1 (see Remark 1). The experimental results are presented in Figures 4 and 5. As shown in Figure 4, the traditional ESC method based on first-order Lie bracket from Dürr et al. (2013) fails to converge to the true minimum of the objective function even after 1200 s, instead settling near (x,y)=(0.98,1.93)(x,y)=(0.98,-1.93). Due to the fourth-order objective function, the traditional ESC method will not be able to excite high-order Lie brackets and hence it is not possible to guarantee exponential convergence for this approach. Instead, the traditional ESC method will follow a polynomial decay rate, which will be much slower than an exponential decay rate as discussed and clarified in Grushkovskaya and Eisa (2025). If given a sufficiently long period of time, then the traditional ESC method in Figure 4 will eventually converge, as noted in the introduction and discussed in Grushkovskaya and Eisa (2025). The comparison in Figure 4 is therefore intended to highlight differences in convergence rates rather than ultimate convergence, with exponential convergence enabling substantially faster practical stabilization within experimentally relevant time horizons.
@@
In contrast, the exponentially convergent unicycle ESC design converges much faster, first reaching the minimum at approximately 350 s and remaining about the true extremum for the duration of the experiment. The new ESC design clearly outperforms the classic literature ESC design and is validated by its ability to reach the minimum in significantly less time than the traditional approach. It is also important to note that a fourth-order objective function exhibits relatively flat regions near the minimum compared to a second-order objective function. The exponentially convergent unicycle ESC design can successfully steer the robot through these flat regions, as evident in Figure 6. Moreover, the variations in the objective function values become very small when the robot is close to the extremum point, confirming that steady-state convergence has been achieved. It is also important to highlight that the proposed design performed well even with expected resolution issues from the motion capturing system feedback due to the very small values of the fourth-order objective function, especially near the extremum. This is a positive indication about the robustness of the proposed design. The reader is directed to watch the experiment in our YouTube channel (Palanikumar et al. (2025a)).
@@
Figure 4: Experimental results for the traditional ESC design based on first-order Lie bracket from Dürr et al. (2013) with a fourth-order objective function. (a) $x$ position (b) $y$ position (c) planar trajectory of the robot. Note that the x-position converges to $0.98$ meters (black-dashed), resulting in a 0.02 meter error, and the y-position converges to 1.93-1.93 (black-dashed), which results in a 0.07 meter error. The true minimum is in red.
@@
Figure 5: Experimental results for the exponentially convergent unicycle ESC design with a fourth-order objective function. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (Palanikumar et al. (2025a)).
@@
The results of this experiment are presented in Figure 7. As shown, the robot successfully reaches and oscillates around the location of maximum light intensity. The light sensor measurements decrease as the robot approaches the source, and the mean value of the signal is minimized toward the end of the experiment, despite small oscillations caused by the robot’s continuous motion. These results further confirm that the proposed unicycle ESC design can autonomously steer the robot through an unknown and spatially varying light field, thereby validating its real-time, model-free source-seeking capability. The reader is directed to watch this experiment video in YouTube (Palanikumar et al. (2025b)).
Figure 7: Experimental results for light source-seeking with the exponentially convergent ESC model. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (Palanikumar et al. (2025b)).
4.4 Further Discussion on the Selection of $\varepsilon$ in our Experiments
Here, we briefly discuss the selection of $\varepsilon$ in Table 2 and Table 3. In both cases, tuning is performed by selecting a perturbation frequency $\omega$, where $\omega=\frac{2\pi}{\varepsilon}$, and then solving for the corresponding value of $\varepsilon$, as discussed in subsection 3.3.
For the case of the known objective function, we selected $\omega=21$, which yields $\varepsilon=0.2992$. This choice was guided by experimental observations and prior experience with the robotic platform as in Elgohary et al. (2025); Bajpai et al. (2024). While larger values of $\omega$ are preferable in theory, the selected value avoids violating hardware limitations discussed in subsections 3.3 and 4.9. In practice, we verified that higher values such as $\omega=30$ and $\omega=42$ were also feasible, but $\omega$ values between 20 and 25 were good enough for the excitation and relatively far from the range of $\omega$ values where hardware limitations are more possible to be encountered.
For the light source-seeking experiment, the objective function is unknown, and therefore the admissible range of $\omega$ (or equivalently $\varepsilon$) may differ from the known objective function case. In particular, the minimum $\omega$ threshold, $\omega^{*}$, required for effective averaging cannot be guaranteed to be identical for the two systems. That is, the maximum $\varepsilon$ threshold, $\varepsilon^{*}$, below which $\varepsilon$ must be chosen, cannot be guaranteed to be the same for when the objective function is known compared to when it is unknown since the objective function is the “output” of the system in ESC designs. Although larger values of $\omega$ remain advantageous in theory (Maggia et al. (2020)), we selected $\omega=42$ to remain within the hardware limitations of the robot. This corresponds to $\varepsilon=0.1496$.
@@ 4.7 Response to Different Initial Values in Simulations and Experiment
Furthermore, we extend the distance between the initial and desired positions and show, in simulation, the ability of the exponentially convergent unicycle ESC to reach the desired position from a farther distance. The initial condition for this simulation is taken as, x(0),y(0)=0.3,0.7x(0),\ y(0)=-0.3,\ -0.7 and the result is shown in Figure 13. The unicycle is still able to converge to the desired position even when starting farther away from the desired position.
Figure 13: Simulation result for an initial position that is farther away from the desired position with the exponentially convergent unicycle ESC. x(0),y(0)=0.3,0.7x(0),\ y(0)=-0.3,\ -0.7.
In the experiment, due to the hardware limitations of the TurtleBot3 robot, we were not able to test initial conditions that were farther away from the desired position. Specifically, the maximum linear velocity of the TurtleBot3 limits the admissible amplitude of the velocity perturbation signal. When the objective function value is large (i.e., far from the extremum), the commanded velocity saturates, which suppresses the effective excitation needed to realize the corresponding Lie bracket direction. As a result, initial conditions that are too far from the desired position cannot be reliably explored in experiment.
Overall, the exponentially convergent unicycle ESC does not show dependence on the initial position directions with respect to the extremum in both simulations and experiments. However, a gap between theory/simulations on the one hand and experiments on the other, was observed as the exponentially convergent unicycle has hardware limitations preventing it from converging if the initial condition is set very far from the extremum; a solution for that could be a future work that develops a bounded update rate design upgrade for the exponentially convergent ESC, analogous to the bounded update design upgrade in Scheinker (2017) with respect to the traditional ESC design in Dürr et al. (2013).
@@ 4.8 Comparison between Simulation and Experiment Results
To begin, we note that the unicycle dynamics model does not perfectly capture the true dynamics of the TurtleBot3 robot. Inertial terms are not captured in the unicycle dynamics as the unicycle is treated as a point mass. Also, dissipation terms, such as friction, are not captured within the unicycle dynamics. For these reasons, it is not possible to directly equate the simulation and experiment parameters. That is, Table 1 and Table 2 are not aligned in terms of the parameters, partly due to the differences in the unicycle model and the real-world system of TurtleBot3. To compare the simulation results with its theoretically predicted behavior vs. the experimental result, we aimed to match the initial convergence rate for the states between the two systems (simulation-based and experiment-based) in a similar fashion to what could be found in ESC literature when one compares the convergence rate of two different ESC systems, see for example the comparisons in Elgohary and Eisa (2025a). Analyzing Table 1 and Table 2, we see that at $t=0$ s, $\dot{y}(0)$ is the same for both systems and the only difference between the two systems is in $\dot{x}(0)$ with the simulation-based system possessing a much higher initial convergence rate in $\dot{x}$ compared to the experiment-based system. That is, for the experiment-based system, $\dot{x}=0.2199$, and for the simulation-based system, $\dot{x}(0)=705.724$. We now introduce a new set of parameters, in Table 4, to reduce the initial convergence rate of $\dot{x}$ in the simulation-based system as much as possible to allow for a relatively fairer comparison with the experiment-based system. The experimental parameters are the same as in Table 2. With the new parameters in Table 4 for the simulation-based system, the initial convergence rate for $\dot{x}$ becomes $\dot{x}(0)=5.5382$. Now, the results that compare the experiment-based system and the simulation-based system, which, theoretically, is guaranteed to have an exponential convergence rate, are provided in Figure 14. We observe that the simulation-based system, even with relatively higher initial convergence rate in $\dot{x}$ (i.e., more advantageous initial speed), is slower than the experimental-based system. This shows that the proposed unicycle ESC follows the theoretical predictions in Section 3 and their simulations; however, one need to keep in mind the reported limitation in Subsection 4.6 where the real-world system may not converge for initial conditions that are very far from the extremum.
@@
4.9 Summary of Observed Hardware Limitations
We note that two primary hardware limitations arise from our experiments.
  • Minimum value of $\varepsilon$. Theoretically, smaller values of $\varepsilon$ (equivalently, larger perturbation frequencies $\omega$) are preferable. In practice, however, there is a lower bound on $\varepsilon$ imposed by the physical limitations of the robot. Large values of $\omega$ challenge the ability of the motors to track the commanded velocity input. Despite this limitation, our experiments demonstrate successful convergence for values of $\varepsilon$ that respect the physical frequency constraints of the platform, as shown in Figures 5, 7, and 12. Additional details on the selection of $\varepsilon$ are provided in subsection 4.4. Moreover, since the control loop is executed with a fixed sampling time, the Nyquist frequency also imposes a lower bound on admissible values of $\varepsilon$.
  • Maximum linear velocity of the robot. The maximum linear velocity of the TurtleBot3 imposes a physical saturation limit on the velocity perturbation signal. As a result, the experimental basin of attraction differs from the theoretical one. This effect becomes particularly evident when the initial condition is far from the extremum. While the unicycle model in simulation can converge from distant initial positions (Fig. 13), the physical robot cannot replicate this behavior due to velocity saturation. For fourth-order objective functions, the objective value grows rapidly with distance from the minimum, leading to large commanded velocities that saturate at the robot’s physical limit. In this regime, the effective excitation required to realize the Lie bracket direction is suppressed, and the robot may oscillate or drift. In contrast, for the experimental cases shown in Figures 5 and 12, the initial objective values remain sufficiently small to allow the perturbation signal to excite the correct descent direction. Consequently, the initial conditions in Fig. 13 cannot be replicated experimentally using the current platform.
@@ 5 Conclusion and future work
This paper provided a novel, first-of-its-kind exponentially convergent unicycle design inspired by recent results on higher-order Lie bracket approximations (Grushkovskaya and Eisa (2025); Pokhrel and Eisa (2023b)). We proved exponential stability of the proposed design for objective functions that behave locally as a fourth-degree polynomial. The theoretical results were validated through extensive simulations and, more importantly, through real-world robotic experiments. In simulations, we demonstrated robust performance of the proposed design under different initial conditions as well as in the presence of measurement noise and delays. In experiments, we validated not only the ability of the proposed design to operate with known objective functions and varying initial conditions, but also its capability to operate in a fully model-free setting through a light source-seeking experiment. Furthermore, a case study comparing experimental and simulation results was presented to confirm that the observed experimental behavior follows the theoretically predicted exponential convergence trends. These results demonstrate that exponentially convergent ESC laws based on higher-order Lie bracket approximations can be practically realized on physical robotic platforms, despite hardware limitations and measurement imperfections.
In future work, we aim to generalize the proposed design to objective functions with higher polynomial degrees and to develop modified ESC structures that reduce oscillations and enforce bounded update rates while preserving exponential convergence properties.